OpenStudy (anonymous):

Solve, using linear combination. 5x + 2y = –4 –3x + 2y = 12 A. (–2, 3) B. (4, 0) C. (2, –7) D. (0, –2)

OpenStudy (anonymous):

Multiply one of the equations by -1 -1(5x+2y=-4) -5x-2y=4 then add the two equations -5x-2y=4 -3x+2y=12 = -8x=16 x=-2 then plug into one of the equations( any will do) 5(-2)+2y=-4 -10+2y=-4 2y=6 y=3 A. (-2,3)